Suppose it is known that the weight of a certain population of individuals are approximately normally distributed with a mean 80 kg and a standard deviation of 16 kg. What is the probability that a person picked at random from this group will weight between 60 and 110 kg?



1
Expert's answer
2022-04-25T13:35:12-0400

We have a normal distribution, "\mu=80, \sigma=16."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"z_1=\cfrac{60-80}{16}=-1.25;\\
z_2=\cfrac{110-80}{16}=1.88.\\
P(60<X<110)=\\
=P(-1.25<Z<1.88)=\\
=P(Z<1.88)-P(Z<-1.25)=\\
=0.96995-0.10565=\\
=0.8643 \text{ (from z-table)} ."